# Flux-Activated Resonant Control of a Bosonic Quantum Memory

### v1

@@ VI Parametric modulation
Although not required by the resonant control protocol, our device architecture is also capable of providing broadband parametric flux modulation of the SQUID.
@@
We show this feature experimentally by using flux pumps to activate parametric resonances between the cavity and transmon with tunable strength. Sweeping the pump frequency $\omega_{m}$ around the transmon-cavity detuning of $\Delta/2\pi\approx 85\,\textnormal{MHz}$ for a variable duration results in the continuous exchange of the transmon excitation with the cavity, as shown by the chevron pattern in Fig. 12(a). The resonant oscillations have a period of $\pi/g_{eff}$, where $g_{eff}$ is the parametric coupling strength.
Figure 12: Protected flux-parametric modulation (a) (b) (c) Qubit spectroscopy with the transmon parked at the cavity frequency with a flux pump $\Phi_{e}\,(t)=A\cos{\big(\omega_{m}t\big)}$ for variable AA (converted to current) and $\omega_{m}/2\pi=\textnormal{60 MHz}$. The frequency split is compared to $g_{\textnormal{eff}}=gJ_{0}\Big(\frac{\textnormal{2.4048}}{\textnormal{90 }\mu\textnormal{A}}\times I\Big)$, shown as a golden dashed line, where $I$ is the current running through the coil.
@@
The parametric coupling strength $g_{\textnormal{eff}}$ can be tuned on demand by changing the flux pump amplitude. Fig. 12(b) shows that the resonant oscillations become faster with larger flux pump amplitude (expressed in units of current), corresponding to an increase in coupling strength. The pump amplitude $I_{p}$ is estimated from the room-temperature input current after going through $26\,\textnormal{dB}$ attenuation and treating the flux line as a short. Fitting the oscillations to a cosine function for each pump amplitude indicates a variable $0\leq g_{eff}\leq 6.6\,\textnormal{MHz}$. However, increasing the pump amplitude after the maximum coupling point leads to lower $g_{eff}$. This is consistent with the expected analytical trend $g_{eff}=g\times J_{1}(I_{p}/\nu)$, where $g$ is the static coupling, $J_{1}$ is a Bessel function of the first kind and $\nu$ is proportionality factor [11].
@@
$\nu=\left(\frac{2\pi}{\Phi_{0}}\frac{d\omega_{T}}{d\varphi_{e}}\frac{L}{\omega_{m}}\right)^{-1},$ (11)
@@
We test strong parametric resonances with flux pumps at higher modulation frequency. This is shown with transmon spectroscopy experiments at different parking detunings $\Delta=63~\textnormal{to}~232\,\textnormal{MHz}$ under a continuous flux pump with frequency $\omega_{m}=\Delta$. The pump creates flux sidebands at frequencies $\bar{\omega}_{T}+N\omega_{m}$, where the average transmon frequency $\bar{\omega}_{T}$ approximately coincides with the parking frequency $\omega_{T}$ for low pump amplitudes. In all spectroscopy plots (Fig. 12(c)), the rightmost feature corresponds to the $N=1$ sideband, which coincides with the cavity frequency and presents a frequency split corresponding to their avoided crossing. The next feature to the left corresponds to the transmon average frequency ($N=0$), and the lower N=1N=-1 and N=2N=-2 sidebands can be seen within range of the spectroscopy for some DC biases. All features display the same frequency splitting indicating the strong coupling of the $N=1$ sideband with the cavity for pump frequencies as high as $232\,\textnormal{MHz}$.
VII Quantum state reconstruction
@@
On the other hand, the Wigner measurement that we performed is expressed as $W(\alpha)=\text{tr}(\hat{P}\hat{D}(\alpha)^{\dagger}\rho\hat{D}(\alpha))$ with $\hat{P}=e^{i\hat{a}^{\dagger}\hat{a}\pi}$ being the parity operator and grid-based displacement points $\{\alpha\}$. Let us expressed all the Wigner data in a vector form $\vec{X}$. Consequently, the input state and Wigner data are related linearly, which can be written in a single matrix equation $\vec{X}=M\vec{Y}+\vec{V}$, where $M$ and $\vec{V}$ denote a mapping matrix and a constant vector, which can easily be computed given the choice of displacements $\{\alpha\}$.
The state reconstruction is carried out in a few steps. First, we performed linear inversion (LI) to get the estimated state parameters $\vec{Y}_{\text{est}}$, given the Wigner measurement data $\vec{X}$: $\vec{Y}_{\text{est}}=M^{+}(\vec{X}-\vec{V})$, where $M^{+}=(M^{\dagger}M)^{-1}M^{\dagger}$ is the left Moore-Penrose pseudoinverse of the matrix $M$. Then, we construct the density matrix $\rho_{\text{LI}}$ based on the estimated parameters. As errors are present in experiments, this density matrix might not be physical, i.e., it might contain negative eigenvalues. Finally, given $\rho_{\text{LI}}$, we employ Bayesian inference (explained in more detail below) to obtain a posterior probability distribution that accurately represent experimental data. From this distribution, we can sample density matrices and compute the mean fidelity and its standard deviation.
VII.1 Bayesian inference
@@
To address this issue, we use Bayesian inference, which is a statistically more accurate approach. Through Bayes’ rule, it gives us a posterior distribution, explicitly showing uncertainty, given the measured experimental data ($\rho_{\text{LI}}$). The posterior distribution allows us to sample any functions of $\rho$ (including itself). For exact implementation of Bayesian inference, we follow an efficient protocol proposed in Ref. [37]. When implementing the protocol, we use the following parameters: $\alpha=1$, which establishes a uniform prior as we do not assume initial preference for the density matrix; the variance of the pseudo-likelihood $\sigma=1/N$ with $N=\#\>\cdot\>(D^{2}-1)$ and #\# is the number of experimental repetition; we also generate $2^{10}$ Markov Chain Monte Carlo samples and apply a thinning parameter of $2^{7}$ to mitigate correlations within the sample chain.
For example, we use the procedure to sample a large number of possible physical density matrices $\{\rho_{\text{BI},i}\}$ and compute their fidelities $F_{i}=(\text{tr}(\sqrt{\sqrt{\rho_{\text{BI},i}}\>\rho_{\text{tar}}\>\sqrt{\rho_{\text{BI},i}}}))^{2}$ to a target state $\rho_{\text{tar}}$. The fidelity and its error presented in the main text are the average fidelity and its standard deviation.
VIII Protocol simulation and error budget
@@
First, we simulate the state preparation of states $|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$ and $|\psi_{2}\rangle=\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ using sequential gaussian pulses to each transition. The system Hamiltonian was reconstructed using the experimentally calibrated transitions from Tab. 3. The cavity lifetime $T_{1,cav}$, transmon lifetime $T_{1}$ and pure dephasing $T_{\phi}=\left(\frac{1}{T_{2}}-\frac{1}{2T_{1}}\right)^{-1}$ are implemented with the respective jump operators. The adiabatic decoupling is taken as ideal and implemented with a change of basis operation. The Wigner tomography executed in the dispersive regime is also reproduced in simulation with experimental parameters, namely a dispersive shift $\chi/2\pi=1.62\,\textnormal{MHz}$, a parity-mapping time of $272\,\textnormal{ns}$ and gaussian $\pi/2$ pulses with $\sigma_{\pi/2}$ of $8\,\textnormal{ns}$ and total duration of $4\sigma$. The simulated tomography is then rescaled following the amplitude of the Wigner function of vacuum obtained with consistent simulation parameters. The state fidelity is then estimated from density matrix reconstruction following the procedure described in Sect. VII.
@@
Figure 13: Simulated state preparation fidelities. Simulated state fidelity of $|\psi_{1}\rangle=\frac{1}{\sqrt{3}}\left(|0\rangle+|3\rangle\right)$ as a function of transmon $T_{1}$ and $T_{\phi}$.
@@
The state preparation is simulated for transmon $T_{1}=10\,\mu\textnormal{s}$ and $15\,\mu\textnormal{s}$, and $T_{2}$ in the $\left[1.3,2.8\right]\,\mu\textnormal{s}$ range. The cavity lifetime is fixed at $T_{1,\textnormal{cav}}=500\,\mu\textnormal{s}$. We first identify that the transmon is highly sensitive to changes in dephasing, as shown from the reconstructed fidelity as function of $T_{\phi}$ in Fig. 13: the fidelity of $|\psi_{1}\rangle$ varies in a large range $87.2\,\%,93.8\,\%]$ ($87.3\,\%,93.8\,\%]$) for $T_{1}=10\,\mu\textnormal{s}$ ($15\,\mu\textnormal{s}$). These values also show the protocol is weakly sensitive to variations in transmon $T_{1}$ in this parameter range. Similar conclusions about $T_{1}$ and $T_{2}$ sensitivity can be drawn for state $|\psi_{2}\rangle$. Simulations with $T_{2}=2.7\,\mu\textnormal{s}$ and $T_{1}=15\,\mu\textnormal{s}$ ($T_{\phi}\approx 2.97\,\mu\textnormal{s}$) lead to $93.2\,\%$ and $87.7\,\%$ for states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, respectively, which reproduce well the experimental fidelities of $93\pm 3\,\%$ and $89\pm 2\,\%$.
We calculate the fidelities of $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ while progressively including more sources of error in the simulation (Tab. 4). The results indicate that the dephasing $T_{\phi}$ has the largest impact in state preparation fidelity, followed by coherent errors due to finite pulse length, and transmon $T_{1}$, in that order. From these simulations, we expect that the quality of state preparation can be significantly enhanced even with modest improvements in the transmon $T_{\phi}$.
We simulate the Givens rotations between Fock states $|1\rangle$ and $|4\rangle$ using the same procedure, with the addition of the optimized $|1+\rangle\leftrightarrow|3$-$\rangle$ state transfer drive. The simulated cavity photon distributions are compared with the experimental reconstructed density matrix in Fig. 4c in the main text, showing a close match.
Simulations indicate that the experimental state fidelities after Givens rotations are significantly affected by coherent errors. When neglecting the transmon $T_{1}$ and $T_{2}$, the state transfer $|1+\rangle\to|3$-$\rangle$ shows a fidelity of $95.0\,\%$ when the system is initialized in $|1+\rangle$. For the same drive, the returning $|3-\rangle\to|1+\rangle$ path shows a fidelity of $85.8\,\%$ when the system is initialized in $|3$-$\rangle$. We attribute these imperfections to the mismatch between the ideal Jaynes-Cummings model used in the drive optimization and the experimental Hamiltonian parameters, which leads to a difference on the order of $\approx 1\,\textnormal{MHz}$ in the frequencies of higher transitions. We expect that further improvements in our calibration and optimization methods will render significant improvement in the Givens rotations fidelities.
@@
Table 4: Error budget for state preparation. State fidelity for $|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$ and $|\psi_{2}\rangle=\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ states with progressive addition of error sources.

### v2

@@ VIII Parametric modulation
Although not required for the resonant control protocol, our device architecture is also capable of providing broadband parametric flux modulation of the SQUID.
@@
We show this feature experimentally by using flux pumps to activate parametric resonances between the cavity and transmon with tunable strength. Sweeping the pump frequency $\omega_{m}$ around the transmon-cavity detuning of $\Delta\approx 85\,\textnormal{MHz}$ for a variable duration results in the continuous exchange of the transmon excitation with the cavity, as shown by the chevron pattern in Fig. S8(a). The resonant oscillations have a period of $\pi/g_{\textnormal{eff}}$, where $g_{\textnormal{eff}}$ is the parametric coupling strength.
Figure S8: Protected flux-parametric modulation. (a) Vacuum Rabi oscillations activated by a flux sideband. The transmon-cavity detuning is $\Delta=\textnormal{85 MHz}$. After being excited with a $\pi$ pulse, a flux pump with $\omega_{m}/2\pi\approx\Delta$ creates a sideband that activates the photon exchange between transmon and cavity for a variable interaction time. (b) Resonant vacuum Rabi oscillations as a function of the flux pump amplitude. The speed of the Rabi oscillations can be continuously tuned in the $g_{\textnormal{eff}}/2\pi=\textnormal{0}\div\textnormal{6.57 MHz}$ range. (c) Transmon spectroscopy for the variable values of $\Delta$ under a flux pump $\varphi_{e}\propto\cos\big(2\pi\Delta t\big)$. The flux pump activates the parametric resonance between the transmon and the cavity, as seen from the frequency split that appears on every sideband due to the avoided crossing.
@@
The parametric coupling strength $g_{\textnormal{eff}}$ can be tuned on demand by changing the flux pump amplitude. Fig. S8(b) shows that the resonant oscillations become faster with larger flux pump amplitude (expressed in units of current), corresponding to an increase in coupling strength. The pump amplitude $I_{p}$ is estimated from the room-temperature input current after going through $26\,\textnormal{dB}$ attenuation and treating the flux line as a short. Fitting the oscillations to a cosine function for each pump amplitude indicates a variable $0\leq g_{\textnormal{eff}}\leq 6.57\,\textnormal{MHz}$. However, increasing the pump amplitude beyond the maximum coupling point leads to lower $g_{\textnormal{eff}}$. This is consistent with the expected analytical trend $g_{\textnormal{eff}}=g\times J_{1}(I_{p}/\nu)$, where $g$ is the static coupling, $J_{1}$ is a Bessel function of the first kind and $\nu$ is proportionality factor [12].
@@
$\nu=\left(\frac{2\pi}{\Phi_{0}}\frac{d\omega_{T}}{d\varphi_{e}}\frac{L}{\omega_{m}}\right)^{-1},$ (S8)
@@
We test strong parametric resonances with flux pumps at higher modulation frequency. This is shown with transmon spectroscopy experiments at different parking detunings $\Delta=63~\textnormal{to}~232\,\textnormal{MHz}$ under a continuous flux pump with frequency $\omega_{m}=\Delta$. The pump creates flux sidebands at frequencies $\bar{\omega}_{T}+N\omega_{m}$, where the average transmon frequency $\bar{\omega}_{T}$ approximately coincides with the parking frequency $\omega_{T}$ for low pump amplitudes. In all spectroscopy plots (Fig. S8(c)), the rightmost feature corresponds to the $N=1$ sideband, which coincides with the cavity frequency and presents a frequency split corresponding to their avoided crossing. The next feature to the left corresponds to the transmon average frequency ($N=0$), and the lower N=1N=-1 and N=2N=-2 sidebands can be seen within range of the spectroscopy for some DC biases. All features display the same frequency splitting indicating the strong coupling of the $N=1$ sideband to the cavity for pump frequencies as high as $232\,\textnormal{MHz}$.
IX Quantum state reconstruction
@@
On the other hand, the Wigner measurement that we performed is expressed as $W(\alpha)=\text{tr}(\hat{P}\hat{D}(\alpha)^{\dagger}\rho\hat{D}(\alpha))$ with $\hat{P}=e^{i\hat{a}^{\dagger}\hat{a}\pi}$ being the parity operator and grid-based displacement points $\{\alpha\}$. Let us express all the Wigner data in a vector form $\vec{X}$. Consequently, the input state and Wigner data are related linearly, which can be written in a single matrix equation $\vec{X}=M\vec{Y}+\vec{V}$, where $M$ and $\vec{V}$ denote a mapping matrix and a constant vector, which can easily be computed given the choice of displacements $\{\alpha\}$.
The state reconstruction is carried out in a few steps. First, we perform linear inversion (LI) to get the estimated state parameters $\vec{Y}_{\text{est}}$, given the Wigner measurement data $\vec{X}$: $\vec{Y}_{\text{est}}=M^{+}(\vec{X}-\vec{V})$, where $M^{+}=(M^{\dagger}M)^{-1}M^{\dagger}$ is the left Moore-Penrose pseudoinverse of the matrix $M$. Then, we construct the density matrix $\rho_{\text{LI}}$ based on the estimated parameters. As errors are present in experiments, this density matrix might not be physical, i.e., it might contain negative eigenvalues. Thus, given $\rho_{\text{LI}}$, we employ Bayesian inference (details below) to obtain a posterior probability distribution that accurately represents experimental data. From this distribution, we can sample density matrices and compute the mean fidelity and its standard deviation.
IX.1 Bayesian inference
@@
To address this issue, we use Bayesian inference, which is a statistically more accurate approach. Through Bayes’ rule, it gives us a posterior distribution, explicitly showing uncertainty, given the measured experimental data ($\rho_{\text{LI}}$). The posterior distribution allows us to sample any functions of $\rho$ (including itself). For exact implementation of Bayesian inference, we follow an efficient protocol proposed in Ref. [46]. When implementing the protocol, we use the following parameters: $\alpha=1$, which establishes a uniform prior as we do not assume initial preference for the density matrix; the variance of the pseudo-likelihood $\sigma=1/N$ with $N=\#\>\cdot\>(D^{2}-1)$, where #\# is the number of experimental repetition; we also generate $2^{10}$ Markov Chain Monte Carlo samples and apply a thinning parameter of $2^{7}$ to mitigate correlations within the sample chain.
For example, we use this procedure to sample a large number of possible physical density matrices $\{\rho_{\text{BI},i}\}$ and compute their fidelities $F_{i}=(\text{tr}(\sqrt{\sqrt{\rho_{\text{BI},i}}\>\rho_{\text{tar}}\>\sqrt{\rho_{\text{BI},i}}}))^{2}$ to a target state $\rho_{\text{tar}}$. The fidelity and its error presented in the main text are the average fidelity and its standard deviation.
The quantum state reconstruction and Bayesian inference routines used in that work are available on GitHub, github.com/Qcrew/Valadares-Dorogov-FAR, along with the raw experimental data (averaged).
X Protocol simulation and error budget
@@
First, we simulate the state preparation of states $|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$ and $|\psi_{2}\rangle=\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ using sequential gaussian pulses for each transition. The system Hamiltonian was reconstructed using the experimentally calibrated transitions from Tab. III. The cavity lifetime $T_{1,\textnormal{cav}}$, transmon lifetime $T_{1}$ and pure dephasing $T_{\phi}=\left(\frac{1}{T_{2}}-\frac{1}{2T_{1}}\right)^{-1}$ are implemented with the respective jump operators. The adiabatic decoupling is taken as ideal and implemented with a change of basis operation. The Wigner tomography executed in the dispersive regime is also reproduced in simulation with experimental parameters, namely a dispersive shift $\chi/2\pi=1.62\,\textnormal{MHz}$, a parity-mapping time of $272\,\textnormal{ns}$ and gaussian $\pi/2$ pulses with $\sigma_{\pi/2}$ of $8\,\textnormal{ns}$ and total duration of $4\sigma$. The simulated tomography is then rescaled following the amplitude of the Wigner function of vacuum obtained with consistent simulation parameters. The state fidelity is then estimated from density matrix reconstruction following the procedure described in Sec. IX.
@@
Figure S9: Simulated state preparation fidelities. Simulated state fidelity of $|\psi_{1}\rangle=\frac{1}{\sqrt{3}}\left(|0\rangle+|3\rangle\right)$ as a function of transmon $T_{1}$ and $T_{\phi}$.
@@
The state preparation is simulated for transmon $T_{1}=10\,\mu\textnormal{s}$ and $15\,\mu\textnormal{s}$, and $T_{2}$ in the $\left[1.3,2.8\right]\,\mu\textnormal{s}$ range. The cavity lifetime is fixed at $T_{1,\textnormal{cav}}=500\,\mu\textnormal{s}$. We first identify that the transmon is highly sensitive to changes in dephasing, as seen from the reconstructed fidelity as function of $T_{\phi}$ in Fig. S9: the fidelity of $|\psi_{1}\rangle$ varies in a large range $87.2\,\%,93.8\,\%]$ ($87.3\,\%,93.8\,\%]$) for $T_{1}=10\,\mu\textnormal{s}$ ($15\,\mu\textnormal{s}$). These values also show the protocol is weakly sensitive to variations in transmon $T_{1}$ in this parameter range. Similar conclusions about $T_{1}$ and $T_{2}$ sensitivity can be drawn for state $|\psi_{2}\rangle$. Simulations with $T_{2}=2.7\,\mu\textnormal{s}$ and $T_{1}=15\,\mu\textnormal{s}$ ($T_{\phi}\approx 2.97\,\mu\textnormal{s}$) lead to $93.2\,\%$ and $87.7\,\%$ for states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, respectively, reproducing well the experimental fidelities of $93\pm 3\,\%$ and $89\pm 2\,\%$.
@@
Table IV: Error budget for state preparation. State fidelity for $|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+|3\rangle\right)$ and $|\psi_{2}\rangle=\tfrac{1}{2}|0\rangle+\tfrac{\sqrt{3}}{\sqrt{8}}\left(i|2\rangle+|4\rangle\right)$ states with progressive addition of error sources.
We calculate the fidelities of $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ while progressively including more sources of error in the simulation (Tab. IV). The results indicate that the dephasing $T_{\phi}$ has the largest impact on state preparation fidelity, followed by coherent errors due to finite pulse length, and transmon $T_{1}$, in that order. From these simulations, we expect that the quality of state preparation can be significantly enhanced even with modest improvements in the transmon $T_{\phi}$.
We simulate the Givens rotations between Fock states $|1\rangle$ and $|4\rangle$ using the same procedure, with the addition of the optimised $|1+\rangle\leftrightarrow|3$-$\rangle$ state transfer drive. The simulated cavity photon distributions are compared with the experimental reconstructed density matrix in Fig. 4c in the main text, showing a close match.
Simulations indicate that the experimental state fidelities after Givens rotations are significantly affected by coherent errors. When neglecting the transmon $T_{1}$ and $T_{2}$, the state transfer $|1+\rangle\to|3$-$\rangle$ shows a fidelity of $95.0\,\%$ when the system is initialised in $|1+\rangle$. For the same drive, the returning $|3-\rangle\to|1+\rangle$ path shows a fidelity of $85.8\,\%$ when the system is initialised in $|3$-$\rangle$. We attribute these imperfections to the mismatch between the ideal Jaynes-Cummings model used in the drive optimization and the experimental Hamiltonian parameters, which leads to a difference on the order of $\approx 1\,\textnormal{MHz}$ in the frequencies of higher transitions. We expect that further improvements in our calibration and optimization methods will render significant improvement in the Givens rotations fidelities.